OpenStudy (anonymous):

The average of 2 and 4 is 3. What is the average of the reciprocals of 2 and 4?

OpenStudy (anonymous):

Take the reciprocal of each number, and add them together. Then, divide the last result by 2 (since there is two numbers).

OpenStudy (anonymous):

1/2 + 1/4 = 2/4 + 1/4 =3/4 3/4 divided by 2= 3/8 3/8 is the answer